Factor x^3-8
I got the answer (x-2)(x^+2x+4) but I don't know how
Please show me step by step to get the answer. Thanks

Answer by confusedgirl81(14)

First this is to understand Quadratic is all bout the finding the zero's.
so x^3 - 8 = 0
x^3 = 8
x = 2
We can use The Difference of Two Cubes Formula to solve this.
The formula says a^3 - b^3. If we factor (a - b) out we will get (a - b)(a^2 + ab + b^2).
So lets put your info in the formula.
let a = x and b = 2
x^3 - 2^3

Now factor x-2 out and we are left with (x-2)(x^2 + 2x + 2^2)
So x^3 - 8 = x^3 - 2^3 = (x-2)(x^2 + 2x + 2^2)


If you need more help, just ask!